# Self-intersection Number of Negative Curves on Fermat Surfaces

### v1

@@ 5. Discussions and Problems
by Theorem 4.5.
The first key obstacle for us to get the Bounded Negativity comes from the term
@@
which may have order $e^{2}$ for $e$ sufficiently large. Moreover, this problem arises from two possible causes: (i) $D\cap E$ contains many singular points of $E$; (ii) $\mu_{p}(E)$ is large for $p\in D\cap E$.
@@
We first consider the case where $D\cap E$ contains as many points as possible. In this case, #$\#(D\cap E)=de$. Then for each $p\in E$, we have $\mu_{p}(E)=0$ by [6, Cor.2.6.8]. Therefore,
@@
5.2. The case $\mu_{p}(E)$ is large for some $p\in D\cap E$.
@@
Does there exist a reduced and irreducible homogeneous polynomial $g_{e_{i}}$ for a sequence $e_{i}\to\infty$ satisfying the following conditions?
@@
    The projective plane curve $V(g_{e_{i}})$ is a rational cuspidal curve with only one singular point which lies on the curve $x^{d}+y^{d}+z^{d}=0$; and
@@
The the answer is positive, then taking an irreducible component of $\mathscr{C}_{e_{i}}$ would disprove the Bounded Negativity Conjecture for $X_{d}$.
5.4. The case $C$ is a smooth rational curve.
If $C$ is a rational curve, then necessarily $k=d$ no matter whether $d$ is prime or not, since in this case, the finite morphism $\rho:C\to D$ reduces to a birational morphism. Therefore, we have
To remove the influence caused by the terms $\mu_{[x,y,z,0]}(C)$, we assume in addition that $C$ is smooth. Then
$C^{2}=e^{2}-(d-1)e-\sum_{p\in D\cap E}\mu_{p}(E).$
Hence we propose the following problem:
Problem 5.5.
Does there exist a reduced and irreducible homogeneous polynomial $g_{e_{i}}$ for a sequence $e_{i}\to\infty$ satisfying the following conditions?
  1. (i)
    The projective plane curve $V(g_{e_{i}})$ is a rational curve; and
  2. (ii)
    The curve $\mathscr{C}_{e_{i}}$ in $\mathbb{P}^{3}$ defined by $g_{e_{i}}(x,y,z)=0,x^{d}+y^{d}+z^{d}+w^{d}=0$ is a reducible curve with smooth rational irreducible components; and
  3. (iii)
    The sequence
    $e_{i}^{2}-(d-1)e_{i}-\sum_{p\in D\cap V(g_{e_{i}})}\mu_{p}(V(g_{e_{i}}))$
    is not bounded from below, where $D$ is the projective plane curve defined by $x^{d}+y^{d}+z^{d}=0$.
The the answer is positive, then the Bounded Negativity Conjecture will not hold for $X_{d}$.
@@ 5.6. An invariant
where $V(g)$ is the projective plane curve defined by $g$.
Problem 5.7.
@@
    More generally, if $f$ defines a smooth projective surface in $\mathbb{P}^{2}$, does it hold that $I_{f}<1$?
@@
    Does the invariant $I_{f}$ depends on the equation of $f$, or at least the degree $d$ of $f$?
@@ 6.1. Puiseux series
The field of fractionary power series, denoted by $\mathbb{C}\langle\langle u\rangle\rangle$ is the direct limit of the fields $\mathbb{C}((u^{\frac{1}{n}}))$, where the morphisms in the directed system are
$\mathbb{C}((u^{\frac{1}{n}}))\to\mathbb{C}((u^{\frac{1}{n^{\prime}}})),\quad n|n^{\prime}\mathrm{\ and\ }\sum a_{i}u^{\frac{i}{n}}\mapsto\sum a_{i}x^{\frac{in^{\prime}/n}{n^{\prime}}}.$
@@
Fractionary power series $s$ with $o_{u}(s)>0$ will be called Puiseux series. After a simultaneous reduction of all the fractional exponents effectively appearing in $s$, we may assume that $n$ and $\gcd\{i:a_{i}\neq 0\}$ are comprime, and then we call $n$ the polydromy order of $s$, denoted by $\nu(s)$. Moreover, we say that $s$ is a convergent fractionary power series if the formal power series $s(t^{n})$ has a positive convergence radius.
@@
Some basic results that we will use in the theory of Puiseux series are summarized as follows; we refer to [6, Chap.1] for the proof.
@@
    $\prod_{\alpha=1}^{b}\left(v-s(e^{2\pi\sqrt{-1}\alpha/b}u)\right)\mathrm{\ divides\ }f\mathrm{\ in\ }\mathbb{C}\{u,v\}.$
@@
    $f=c(u,v)\prod_{\alpha=1}^{b}\left(v-s(e^{2\pi\sqrt{-1}\alpha/b}u)\right)$
@@ 6.3. Resultants
$\mathrm{Res}\,(f,g)=a_{0}^{e}b_{0}^{d}\prod_{\stackrel{{\scriptstyle 1\leq i\leq d}}{{1\leq j\leq e}}}(\lambda_{i}-\mu_{j}).$
@@
    $\mathrm{Disc}\,(f)=(-1)^{d(d-1)/2}a_{0}^{2d-1}\prod_{i\neq j}(\lambda_{i}-\lambda_{j})=\frac{(-1)^{d(d-1)/2}}{a_{0}}\mathrm{Res}\,\left(f,\frac{\mathrm{d}f}{\mathrm{d}v}\right).$
@@
    $\mathrm{Disc}\,(f)=\left(\prod_{i=1}^{k}\mathrm{Disc}\,(f)\right)\left(\prod_{1\leq i<j\leq k}\mathrm{Res}\,(f_{i},f_{j})^{2}\right).$
@@ 6.5. Intersection multiplicities
Now we use the Puiseux series and resultants to study intersection multiplicities of plane curve germs. Suppose $F,G$ are two reduced curve germs in $(\mathbb{C}^{2},0)$, with isolated singularity at 00, defined by
@@
where $a_{i},b_{j}\in\mathbb{C}\{u\}$ satisfy $a_{i}(0)=b_{j}(0)=0$ for $i\leq d,j\leq e$. For convenience, we will also regard $a_{i},b_{j}$ as elements of $\mathbb{C}\langle\langle u\rangle\rangle$ by the inclusion $\mathbb{C}\{u\}\subset\mathbb{C}\langle\langle u\rangle\rangle$.
@@
    $i_{0}(f,g)=do_{u}\left(g(u,s(u))\right)=\sum_{\alpha=1}^{d}o_{u}\left(g\left(u,s(e^{2\pi\sqrt{-1}\alpha/d}u)\right)\right).$
@@
    $i_{0}(f,g)=\sum_{\stackrel{{\scriptstyle 1\leq i\leq d}}{{1\leq j\leq e}}}o_{u}(s_{i}-s_{j}^{\prime}).$
@@
To give the proof of Proposition 4.4, we first recall our settings. We are given a reduced and irreducible curve $C$ on $X_{d}$, from which we get the following objects:
@@
    Fix an intersection point $P=[x_{0},y_{0},z_{0},w_{0}]$ of different irreducible components of $\mathscr{C}$ satisfying $w_{0}=0$. We assume $x_{0}\neq 0$ and $z_{0}\neq 0$, and denote $p=[x_{0},y_{0},z_{0}]$.
@@
where $h_{1},\cdots,h_{r}$ are irreducible Weierstrass polynomial in $\mathbb{C}\{u\}[v]$ of order $b_{1},\cdots,b_{r}$ respectively, and $h_{0}$ is a nonzero constant. Recall that by assumption $h$ is $v$-general of order $b$, hence $b=b_{1}+\cdots+b_{r}$.
@@
$h_{i}=\prod_{\alpha=1}^{b_{i}}\left(v-s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right),\ i\leq r.$
@@
$\phi_{i}(u,\omega)=\prod_{\alpha=1}^{b_{i}}\left(\omega^{d}+l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right)\right),\ i\leq r,$
@@
Then $\Phi$ is a Weierstrass polynomial in $\omega$ of order $bd$ and moreover, the scheme $\mathscr{C}$ is locally defined by the equation $\Phi(u,\omega)=0$ in the local coordinate system $(u,\omega)$ on $X_{d}$.
@@
$o_{u}\left(l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right)\right)=\frac{i_{0}(h_{i},l)}{b_{i}}\leq\frac{i_{p}(D,E)}{b_{i}}<\infty.$
Therefore, $l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right)\neq 0$ and thus $\omega^{d}+l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right)$ is reduced. Moreover, recall that
@@
$l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right)-l\left(u,s_{i}(e^{2\pi\sqrt{-1}\beta/b_{i}}u)\right)$
$=$ $\left(s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)-s_{i}(e^{2\pi\sqrt{-1}\beta/b_{i}}u)\right)\left(d\left(\frac{z_{0}}{x_{0}}\right)^{d-1}+r(u)\right)$
$\simeq$ $s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)-s_{i}(e^{2\pi\sqrt{-1}\beta/b_{i}}u)$
@@
$s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)-s_{i}(e^{2\pi\sqrt{-1}\beta/b_{i}}u)\neq 0,$
@@
$\gcd\left(\omega^{d}+l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right),\omega^{d}+l\left(u,s_{i}(e^{2\pi\sqrt{-1}\beta/b_{i}}u)\right)\right)=1.$
@@
$l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right)\neq l\left(u,s_{j}(e^{2\pi\sqrt{-1}\beta/b_{j}}u)\right).$
@@
$l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right)-l\left(u,s_{j}(e^{2\pi\sqrt{-1}\beta/b_{j}}u)\right)\simeq s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)-s_{j}(e^{2\pi\sqrt{-1}\beta/b_{j}}u).$
Recall that $h_{i}$ and $h_{j}$ are coprime, they do not have common roots in $\mathbb{C}\langle\langle u\rangle\rangle$. Therefore, $s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)-s_{j}(e^{2\pi\sqrt{-1}\beta/b_{j}}u)\neq 0$.
@@
$=$ $\sum_{\alpha=1}^{b_{i}}o_{u}\left(\mathrm{Disc}\,\left(\omega^{d}+l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right)\right)\right)$
$+\sum_{1\leq\alpha<\beta\leq b_{i}}o_{u}\left(\mathrm{Res}\,\left(\omega^{d}+l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right),\omega^{d}+l\left(u,s_{i}(e^{2\pi\sqrt{-1}\beta/b_{i}}u)\right)\right)^{2}\right).$
@@
    $\sum_{\alpha=1}^{b_{i}}o_{u}\left(\mathrm{Disc}\,\left(\omega^{d}+l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right)\right)\right)=(d-1)i_{0}(h_{i},l),$
@@
    $\sum_{1\leq\alpha<\beta\leq b_{i}}o_{u}\left(\mathrm{Res}\,\left(\omega^{d}+l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right),\omega^{d}+l\left(u,s_{i}(e^{2\pi\sqrt{-1}\beta/b_{i}}u)\right)\right)^{2}\right)$
@@
$\mathrm{Disc}\,\left(\omega^{d}+l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right)\right)$ $\simeq$ $\mathrm{Res}\,\left(\omega^{d}+l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right),w^{d-1}\right)$
$\simeq$ $\left(l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right)\right)^{d-1};$
@@
$o_{u}\left(\mathrm{Disc}\,\left(\omega^{d}+l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right)\right)\right)$ $=$ $(d-1)\left(l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right)\right)$
@@
Summing over $\alpha=1,\cdots,b_{i}$ gives the desired equality.
@@
$\mathrm{Res}\,\left(\omega^{d}+l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right),\omega^{d}+l\left(u,s_{i}(e^{2\pi\sqrt{-1}\beta/b_{i}}u)\right)\right)$
$=$ $\mathrm{Res}\,\left(\omega^{d}+l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right),l\left(u,s_{i}(e^{2\pi\sqrt{-1}\beta/b_{i}}u)\right)-l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right)\right)$
$\simeq$ $\left(l\left(u,s_{i}(e^{2\pi\sqrt{-1}\beta/b_{i}}u)\right)-l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right)\right)^{d}.$
@@
$l\left(u,s_{i}(e^{2\pi\sqrt{-1}\beta/b_{i}}u)\right)-l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right)\simeq s_{i}(e^{2\pi\sqrt{-1}\beta/b_{i}}u)-s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u).$
@@
$\mathrm{Res}\,\left(\omega^{d}+l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right),\omega^{d}+l\left(u,s_{i}(e^{2\pi\sqrt{-1}\beta/b_{i}}u)\right)\right)$
$\simeq$ $\left(s_{i}(e^{2\pi\sqrt{-1}\beta/b_{i}}u)-s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right)^{d}.$
@@
$\sum_{1\leq\alpha<\beta\leq b_{i}}o_{u}\left(\mathrm{Res}\,\left(\omega^{d}+l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right),\omega^{d}+l\left(u,s_{i}(e^{2\pi\sqrt{-1}\beta/b_{i}}u)\right)\right)^{2}\right)$
$=$ $d\sum_{1\leq\alpha<\beta\leq b_{i}}o_{u}\left(s_{i}(e^{2\pi\sqrt{-1}\beta/b_{i}}u)-s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right)^{2}$
@@
$=$ $\prod_{\alpha=1}^{b_{i}}\prod_{\beta=1}^{b_{j}}\mathrm{Res}\,\left(\omega^{d}+l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right),\omega^{d}+l\left(u,s_{j}(e^{2\pi\sqrt{-1}\beta/b_{j}}u)\right)\right)^{2}$
$\simeq$ $\prod_{\alpha=1}^{b_{i}}\prod_{\beta=1}^{b_{j}}\mathrm{Res}\,\left(\omega^{d}+l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right),l\left(u,s_{j}(e^{2\pi\sqrt{-1}\beta/b_{j}}u)\right)-l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right)\right)^{2}$
$\simeq$ $\prod_{\alpha=1}^{b_{i}}\prod_{\beta=1}^{b_{j}}\left(l\left(u,s_{j}(e^{2\pi\sqrt{-1}\beta/b_{j}}u)\right)-l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right)\right)^{2d}.$
@@
$l\left(u,s_{j}(e^{2\pi\sqrt{-1}\beta/b_{j}}u)\right)-l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right)\simeq s_{j}(e^{2\pi\sqrt{-1}\beta/b_{j}}u)-s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u);$
@@
$\simeq$ $\prod_{\alpha=1}^{b_{i}}\prod_{\beta=1}^{b_{j}}\left(s_{j}(e^{2\pi\sqrt{-1}\beta/b_{j}}u)-s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right)^{2d}$
@@
(C) Since $C_{1},\cdots,C_{k}$ are the irreducible components of $\mathscr{C}$, we have
@@
Note that $P$ is fixed under the Galois group for the branched cover $\rho$, and the Galois group acts transitively on $\{C_{1},\cdots,C_{k}\}$, hence
$\mu_{P}(C_{i})=\mu_{P}(C),\ \mathrm{for\ all\ }i=1,\cdots,k.$

### v2

@@ 5. Discussions and Problems
This equality is from Theorem 4.5.
The first key obstacle for us to prove the Bounded Negativity Conjecture on $X_{d}$ comes from the term
@@
which may be comparable to $e^{2}-(d-1)e$ for $e$ sufficiently large. Moreover, this problem arises from two possible causes: (i) $D\cap E$ contains many singular points of $E$; (ii) $\mu_{p}(E)$ is large for $p\in D\cap E$.
@@
We first consider the case where $D\cap E$ contains as many points as possible, i.e., #$\#(D\cap E)=de$. Then for each $p\in D\cap E$, $D$ and $E$ intersect transversely at $p$ by [6, Cor.2.6.8]. In particular, $p$ is a smooth point of $E$, and hence $\mu_{p}(E)=0$. Therefore,
@@
5.2. The case where $\mu_{p}(E)$ is large for some $p\in D\cap E$.
@@
Does there exist a sequence of irreducible homogeneous polynomials $g_{e_{i}}$ of degrees $e_{i}$ for a sequence of integers $e_{i}\to\infty$ such that the following conditions are satisfied?
@@
    The projective plane curve $V(g_{e_{i}})$ is a rational cuspidal curve with exactly one singular point lying on the curve $x^{d}+y^{d}+z^{d}=0$; and
@@
If the answer is positive, then taking an irreducible component $C_{e_{i}}$ of $\mathscr{C}_{e_{i}}$, we will have
$C_{e_{i}}^{2}=\frac{d}{k_{i}}\left((4-d)e_{i}-1\right)-1\to-\infty$
where $2\leq k_{i}\leq d$. This would disprove the Bounded Negativity Conjecture for $X_{d}$.
@@ 5.4. An invariant
where $V(g)$ denotes the projective plane curve defined by $g$.
Problem 5.5.
@@
    More generally, if $f$ defines a smooth projective curve in $\mathbb{P}^{2}$, does it hold that $I_{f}<1$?
@@
    Does the invariant $I_{f}$ depend on the equation of $f$, or at least the degree $d$ of $f$?
@@ 6.1. Puiseux series
The field of fractionary power series, denoted by $\mathbb{C}\langle\langle u\rangle\rangle$ is the direct limit of the fields $\mathbb{C}((u^{\frac{1}{n}}))$, where the morphisms in the directed system are given by
$\mathbb{C}((u^{\frac{1}{n}}))\to\mathbb{C}((u^{\frac{1}{n^{\prime}}})),\quad n|n^{\prime}\mathrm{\ and\ }\sum a_{i}u^{\frac{i}{n}}\mapsto\sum a_{i}u^{\frac{in^{\prime}/n}{n^{\prime}}}.$
@@
Fractionary power series $s$ with $o_{u}(s)>0$ will be called Puiseux series. After a simultaneous reduction of all the fractional exponents effectively appearing in $s$, we may assume that $n$ and $\gcd\{i:a_{i}\neq 0\}$ are coprime, and then we call $n$ the polydromy order of $s$, denoted by $\nu(s)$. We will also use the following notation: if $n=\nu(s)$ and $s$ can be written as
$s(u)=\sum a_{i}u^{\frac{i}{n}},$
then define
$s_{,\alpha}(u)=\sum e^{2\pi\sqrt{-1}\alpha i/n}a_{i}u^{\frac{i}{n}}$
for $\alpha=1,\ldots,n$. Moreover, we say that $s$ is a convergent fractionary power series if the formal power series $s(t^{n})$ has a positive convergence radius.
@@
Some basic results which we will implicitly use in the theory of Puiseux series are summarized as follows; we refer to [6, Chap.1] for the proof.
@@
    $\prod_{\alpha=1}^{b}\left(v-s_{,\alpha}(u)\right)\mathrm{\ divides\ }f\mathrm{\ in\ }\mathbb{C}\{u,v\}.$
@@
    $f=c(u,v)\prod_{\alpha=1}^{b}\left(v-s_{,\alpha}(u)\right)$
@@ 6.3. Resultants
$\mathrm{Res}(f,g)=a_{0}^{e}b_{0}^{d}\prod_{\begin{subarray}{c}1\leq i\leq d\\ 1\leq j\leq e\end{subarray}}(\lambda_{i}-\mu_{j}).$
@@
    $\mathrm{Disc}(f)=(-1)^{d(d-1)/2}a_{0}^{2d-2}\prod_{i\neq j}(\lambda_{i}-\lambda_{j})=\frac{(-1)^{d(d-1)/2}}{a_{0}}\mathrm{Res}\left(f,\frac{\partial f}{\partial v}\right).$
@@
    $\mathrm{Disc}(f)=\left(\prod_{i=1}^{k}\mathrm{Disc}(f_{i})\right)\left(\prod_{1\leq i<j\leq k}\mathrm{Res}(f_{i},f_{j})^{2}\right).$
@@ 6.5. Intersection multiplicities
Now we use the Puiseux series and resultants to study intersection multiplicities of plane curve germs. Suppose $F,G$ are two reduced curve germs in $(\mathbb{C}^{2},0)$, defined by
@@
where $a_{i},b_{j}\in\mathbb{C}\{u\}$ satisfy $a_{i}(0)=b_{j}(0)=0$ for $1\leq i\leq d,1\leq j\leq e$. For convenience, we will also regard $a_{i},b_{j}$ as elements of $\mathbb{C}\langle\langle u\rangle\rangle$ by the inclusion $\mathbb{C}\{u\}\subset\mathbb{C}\langle\langle u\rangle\rangle$.
@@
    $i_{0}(f,g)=do_{u}\left(g(u,s(u))\right)=\sum_{\alpha=1}^{d}o_{u}\left(g\left(u,s_{,\alpha}(u)\right)\right).$
@@
    $i_{0}(f,g)=\sum_{\begin{subarray}{c}1\leq i\leq d\\ 1\leq j\leq e\end{subarray}}o_{u}(s_{i}-s_{j}^{\prime}).$
@@
Before presenting the proof of Proposition 4.4, we first recall our settings. We are given a reduced and irreducible curve $C$ on $X_{d}$, from which we obtain the following objects:
@@
    Fix an intersection point $P=[x_{0},y_{0},z_{0},w_{0}]$ of different irreducible components of $\mathscr{C}$. We have $w_{0}=0$. We also assume $x_{0}\neq 0$ and $z_{0}\neq 0$, and denote $p=[x_{0},y_{0},z_{0}]$.
@@
where $h_{1},\ldots,h_{r}$ are irreducible Weierstrass polynomial in $\mathbb{C}\{u\}[v]$ of degrees $b_{1},\ldots,b_{r}$ respectively, and $h_{0}$ is an invertible element in $\mathbb{C}\{u,v\}$. Recall that by assumption $h$ is $v$-general of order $b$, hence $b=b_{1}+\cdots+b_{r}$.
@@
$h_{i}=\prod_{\alpha=1}^{b_{i}}\left(v-s_{i,\alpha}(u)\right),\ i\leq r.$
(Recall that by our convention, if $s_{i}$ is written as
$s_{i}(u)=\sum a_{k}u^{\frac{k}{b_{i}}}$
then
$s_{i,\alpha}(u)=\sum e^{2\pi\sqrt{-1}\alpha k/b_{i}}a_{k}u^{\frac{k}{b_{i}}}$
for $\alpha=1,\ldots,b_{i}$.)
@@
$\phi_{i}(u,\omega)=\prod_{\alpha=1}^{b_{i}}\left(\omega^{d}+l\left(u,s_{i,\alpha}(u)\right)\right),\ i\leq r,$
@@
Then $\Phi$ is a Weierstrass polynomial in $\mathbb{C}\{u\}[\omega]$ of degree $bd$ and moreover, the scheme $\mathscr{C}$ is locally defined by the equation $\Phi(u,\omega)=0$ in the local coordinate system $(u,\omega)$ on $X_{d}$.
@@
$o_{u}\left(l\left(u,s_{i,\alpha}(u)\right)\right)=\frac{i_{0}(h_{i},l)}{b_{i}}\leq\frac{i_{p}(D,E)}{b_{i}}<\infty.$
In particular, $l\left(u,s_{i,\alpha}(u)\right)\neq 0$; and thus $\omega^{d}+l\left(u,s_{i,\alpha}(u)\right)$ is reduced. Moreover, recall that
@@
$l\left(u,s_{i,\alpha}(u)\right)-l\left(u,s_{i,\beta}(u)\right)$
$=$ $\left(s_{i,\alpha}(u)-s_{i,\beta}(u)\right)\left(d\left(\frac{z_{0}}{x_{0}}\right)^{d-1}+r(u)\right)$
$\simeq$ $s_{i,\alpha}(u)-s_{i,\beta}(u)$
@@
$s_{i,\alpha}(u)-s_{i,\beta}(u)\neq 0,$
@@
$\gcd\left(\omega^{d}+l\left(u,s_{i,\alpha}(u)\right),\omega^{d}+l\left(u,s_{i,\beta}(u)\right)\right)=1.$
@@
$l\left(u,s_{i,\alpha}(u)\right)\neq l\left(u,s_{j,\beta}(u)\right).$
@@
$l\left(u,s_{i,\alpha}(u)\right)-l\left(u,s_{j,\beta}(u)\right)\simeq s_{i,\alpha}(u)-s_{j,\beta}(u).$
Recall that $h_{i}$ and $h_{j}$ are coprime, so they do not have common roots in $\mathbb{C}\langle\langle u\rangle\rangle$. Therefore $s_{i,\alpha}(u)-s_{j,\beta}(u)\neq 0$.
@@
$=$ $\sum_{\alpha=1}^{b_{i}}o_{u}\left(\mathrm{Disc}\left(\omega^{d}+l\left(u,s_{i,\alpha}(u)\right)\right)\right)$
$+\sum_{1\leq\alpha<\beta\leq b_{i}}o_{u}\left(\mathrm{Res}\left(\omega^{d}+l\left(u,s_{i,\alpha}(u)\right),\omega^{d}+l\left(u,s_{i,\beta}(u)\right)\right)^{2}\right).$
@@
    $\sum_{\alpha=1}^{b_{i}}o_{u}\left(\mathrm{Disc}\left(\omega^{d}+l\left(u,s_{i,\alpha}(u)\right)\right)\right)=(d-1)i_{0}(h_{i},l),$
@@
    $\sum_{1\leq\alpha<\beta\leq b_{i}}o_{u}\left(\mathrm{Res}\left(\omega^{d}+l\left(u,s_{i,\alpha}(u)\right),\omega^{d}+l\left(u,s_{i,\beta}(u)\right)\right)^{2}\right)$
@@
$\mathrm{Disc}\left(\omega^{d}+l\left(u,s_{i,\alpha}(u)\right)\right)$ $\simeq$ $\mathrm{Res}\left(\omega^{d}+l\left(u,s_{i,\alpha}(u)\right),\omega^{d-1}\right)$
$\simeq$ $\left(l\left(u,s_{i,\alpha}(u)\right)\right)^{d-1};$
@@
$o_{u}\left(\mathrm{Disc}\left(\omega^{d}+l\left(u,s_{i,\alpha}(u)\right)\right)\right)$ $=$ $(d-1)\left(l\left(u,s_{i,\alpha}(u)\right)\right)$
@@
Summing over $\alpha=1,\ldots,b_{i}$ gives the desired equality.
@@
$\mathrm{Res}\left(\omega^{d}+l\left(u,s_{i,\alpha}(u)\right),\omega^{d}+l\left(u,s_{i,\beta}(u)\right)\right)$
$=$ $\mathrm{Res}\left(\omega^{d}+l\left(u,s_{i,\alpha}(u)\right),l\left(u,s_{i,\beta}(u)\right)-l\left(u,s_{i,\alpha}(u)\right)\right)$
$\simeq$ $\left(l\left(u,s_{i,\beta}(u)\right)-l\left(u,s_{i,\alpha}(u)\right)\right)^{d}.$
@@
$l\left(u,s_{i,\beta}(u)\right)-l\left(u,s_{i,\alpha}(u)\right)\simeq s_{i,\beta}(u)-s_{i,\alpha}(u).$
@@
$\mathrm{Res}\left(\omega^{d}+l\left(u,s_{i,\alpha}(u)\right),\omega^{d}+l\left(u,s_{i,\beta}(u)\right)\right)$
$\simeq$ $\left(s_{i,\beta}(u)-s_{i,\alpha}(u)\right)^{d}.$
@@
$\sum_{1\leq\alpha<\beta\leq b_{i}}o_{u}\left(\mathrm{Res}\left(\omega^{d}+l\left(u,s_{i,\alpha}(u)\right),\omega^{d}+l\left(u,s_{i,\beta}(u)\right)\right)^{2}\right)$
$=$ $d\sum_{1\leq\alpha<\beta\leq b_{i}}o_{u}\left(s_{i,\beta}(u)-s_{i,\alpha}(u)\right)^{2}$
@@
$=$ $\prod_{\alpha=1}^{b_{i}}\prod_{\beta=1}^{b_{j}}\mathrm{Res}\left(\omega^{d}+l\left(u,s_{i,\alpha}(u)\right),\omega^{d}+l\left(u,s_{j,\beta}(u)\right)\right)^{2}$
$\simeq$ $\prod_{\alpha=1}^{b_{i}}\prod_{\beta=1}^{b_{j}}\mathrm{Res}\left(\omega^{d}+l\left(u,s_{i,\alpha}(u)\right),l\left(u,s_{j,\beta}(u)\right)-l\left(u,s_{i,\alpha}(u)\right)\right)^{2}$
$\simeq$ $\prod_{\alpha=1}^{b_{i}}\prod_{\beta=1}^{b_{j}}\left(l\left(u,s_{j,\beta}(u)\right)-l\left(u,s_{i,\alpha}(u)\right)\right)^{2d}.$
@@
$l\left(u,s_{j,\beta}(u)\right)-l\left(u,s_{i,\alpha}(u)\right)\simeq s_{j,\beta}(u)-s_{i,\alpha}(u);$
@@
$\simeq$ $\prod_{\alpha=1}^{b_{i}}\prod_{\beta=1}^{b_{j}}\left(s_{j,\beta}(u)-s_{i,\alpha}(u)\right)^{2d}$
@@
(C) Since $C_{1},\ldots,C_{k}$ are the irreducible components of $\mathscr{C}$, we have
@@
Note that $P$ is a fixed point under the action of the Galois group for the branched cover $\rho$, and the Galois group acts transitively on $\{C_{1},\ldots,C_{k}\}$, hence
$\mu_{P}(C_{i})=\mu_{P}(C),\ \mathrm{for\ all\ }i=1,\ldots,k.$